# Thermodynamic Origin of Degree-Day Scaling in Phase-Change Systems

### v1

@@ Abstract
Phase transitions impose topological constraints on thermodynamic state variables, masking energetic fluctuations at the phase boundary. This constraint is most apparent in melting systems, where temperature remains pinned despite continued energy input. Here we resolve this information loss by introducing a latent temperature—a counterfactual trajectory describing the system’s unconstrained thermal evolution. We show that energy conservation alone enforces a rigorous duality between the total latent heat dissipated during phase change and the accumulated exceedance of the latent temperature above the melting point. This duality is mathematically equivalent to the one-dimensional Wasserstein-1 distance between the latent and observed temperature trajectories, with the transport cost set by a characteristic surface dissipation timescale and melting energy. Applied to ice-sheet surface melting, this timescale admits a direct physical interpretation in terms of radiative and turbulent heat loss. The same framework yields a first-principles derivation of the empirical Positive Degree-Day law and predicts realistic degree-day factors that emerge from surface energy balance, without ad hoc calibration. More broadly, phase change emerges as an optimal transport process that projects continuous energetic variability onto a constrained thermodynamic boundary.
I Introduction
@@
A prominent yet heuristic circumvention of this degeneracy is the Positive Degree-Day (PDD) scaling [5], which remains a standard parameterization in Earth System Models for projecting cryospheric responses across diverse timescales [8, 9, 7]. To estimate melt rates, this empirical approach assumes that cumulative mass loss is proportional to the time-integrated air temperature above the freezing point [1]. While the PDD method exhibits remarkable predictive skill under relatively stable climatic regimes [14], its lack of a first-principles derivation from the surface energy balance introduces structural uncertainties. This limitation becomes particularly acute under non-stationary climate conditions, where the historical coupling between air temperature and the latent heat flux may diverge, potentially rendering existing parameterizations unreliable for future projections [10]. A fundamental gap therefore persists between the rigorous physics of energy-balance formulations and the practical efficacy of temperature-index schemes.
@@ II Latent Temperature
Geometrically, since $\theta_{\mathrm{obs}}(t)=\theta_{f}$ throughout the melt season, the left-hand side of Eq. (6) corresponds to the $L^{1}$ distance between the latent and observed temperature trajectories. In one dimension, this distance is equivalent to the Wasserstein-1 metric. Phase change may therefore be interpreted as an optimal transport process that compresses latent thermal fluctuations onto the melting boundary, with the melt energy determining the associated transport cost (see Supplementary Section S1).
@@ III Physical Interpetation of PDD and Ideal Case
Using representative values for ice-sheet surfaces ($C_{\rm sen}\approx 22.5\,\mathrm{W\,m^{-2}\,K^{-1}}$, $\theta_{f}=273.15\,\mathrm{K}$, $\sigma=5.67\times 10^{-8}\,\mathrm{W\,m^{-2}\,K^{-4}}$, $\rho_{i}=917\,\mathrm{kg\,m^{-3}}$, and $L_{f}=3.34\times 10^{5}\,\mathrm{J\,kg^{-1}}$) [3, 13], we obtain $C_{\rm PDD}\approx 7.6\,\mathrm{mm\,d^{-1}\,{}^{\circ}C^{-1}}$, well within the empirically inferred range [12]. This agreement indicates that the relaxation timescale $\tau$ emerging from the latent framework is quantitatively consistent with known radiative and turbulent dissipation processes.
@@
with amplitude $A=10$. The function is chosen such that the positive forcing phase spans approximately two-thirds of the cycle to better display. For clarity, we adopt a normalized thermal inertia $\rho C_{s}=1$ and reconstruct the latent temperature $\theta(t)$ subject to the energetic constraints described above and minimum $\mathcal{R}$ variation (see Supplementary S2).
@@ Acknowledgements.
Supplementary Materials for Thermodynamic Origin of Degree-Day Scaling in Phase-Change Systems
S1 S1. Latent temperature and the 1-Wasserstein distance
@@
$\int_{0}^{1}\bigl[\theta(t)-\theta_{\mathrm{obs}}(t)\bigr]\,dt=\int_{0}^{1}\left|F_{\theta}^{-1}(p)-F_{\theta_{\mathrm{obs}}}^{-1}(p)\right|\,dp.$ (S2)
@@
S2 S2. Supplementary Methods: Thermodynamic Origin of Degree-Day Scaling in Phase-Change Systems
S2.1 Governing Equation and General Solution
@@
$\rho C_{s}\frac{d\theta}{dt}+\frac{1}{\tau}(\theta-\theta_{f})=Q(t)-\mathcal{R}(t),$ (S3)
@@
Equation (S3) admits the formal solution
$\theta(t)=\theta_{f}+e^{-\lambda(t-t_{s})}\int_{t_{s}}^{t}\frac{Q(s)-\mathcal{R}(s)}{\rho C_{s}}e^{\lambda(s-t_{s})}\,ds,$ (S4)
@@
$\int_{t_{s}}^{t_{e}}\mathcal{R}(t)\,dt=0.$ (S5)
Second, closure of the latent trajectory at the end of the melt season, $\theta(t_{e})=\theta_{f}$, applied to Eq. (S4), yields the weighted constraint
$\int_{t_{s}}^{t_{e}}\mathcal{R}(t)e^{\lambda t}\,dt=\int_{t_{s}}^{t_{e}}Q(t)e^{\lambda t}\,dt\equiv K.$ (S6)
@@
S2.3 Minimum Structural Perturbation Principle
@@
$J[\mathcal{R}]=\int_{t_{s}}^{t_{e}}\mathcal{R}(t)^{2}\,dt,$ (S7)
subject to the constraints (S5) and (S6).
@@
$\mathcal{L}=\int_{t_{s}}^{t_{e}}\mathcal{R}^{2}\,dt+\mu_{1}\int_{t_{s}}^{t_{e}}\mathcal{R}\,dt+\mu_{2}\left(\int_{t_{s}}^{t_{e}}\mathcal{R}e^{\lambda t}\,dt-K\right).$ (S8)
@@
$\frac{\delta\mathcal{L}}{\delta\mathcal{R}}=2\mathcal{R}(t)+\mu_{1}+\mu_{2}e^{\lambda t}=0,$ (S9)
@@
$\mathcal{R}^{*}(t)=\alpha+\beta e^{\lambda t},$ (S10)
@@
S2.4 Determination of Coefficients
Substituting Eq. (S10) into the two admissibility constraints yields the linear system
$\begin{pmatrix}\displaystyle\int_{t_{s}}^{t_{e}}dt&\displaystyle\int_{t_{s}}^{t_{e}}e^{\lambda t}\,dt\\[5.69054pt] \displaystyle\int_{t_{s}}^{t_{e}}e^{\lambda t}\,dt&\displaystyle\int_{t_{s}}^{t_{e}}e^{2\lambda t}\,dt\end{pmatrix}\begin{pmatrix}\alpha\\ \beta\end{pmatrix}=\begin{pmatrix}0\\ K\end{pmatrix}$, (S11)
where $K$ is defined by Eq. (S6). All integrals are analytic, ensuring a smooth and numerically stable solution for $\mathcal{R}^{*}(t)$ and $\theta(t)$ for any prescribed $\tau$.
S2.5 Selection of the Dissipation Timescale
@@
    Solve Eq. (S11) for $\alpha$ and $\beta$.
@@
    Reconstruct $\theta(t)$ from Eq. (S4) using $\mathcal{R}^{*}(t)$.

### v2

@@ Abstract
The temporal evolution of systems undergoing phase changes is constrained by thermodynamic thresholds, pinning state variables to fixed values. This threshold behavior redirects energy exchange into phase transitions, rendering the underlying energetics inaccessible from temperature alone and introducing strong nonlinearity. Here, we introduce the concept of latent temperature—a counterfactual thermal trajectory reconstructed by suppressing phase change, thereby recovering the underlying unconstrained evolution. We show that the latent trajectory is uniquely determined by a variational principle of minimum structural perturbation, minimizing nonlinear residuals relative to a linear relaxation process. The cumulative temperature elevation—mathematically equivalent to the one-dimensional Wasserstein distance between the latent and observed temperature distributions—is strictly proportional to the latent heat of phase change. The proportionality constant emerges naturally as the characteristic dissipation timescale $\tau$. This result provides a first-principles derivation of the classical Positive Degree-Day relationship and establishes phase-change thermodynamics as an optimal transport problem in non-equilibrium systems.
preprint: APS/123-QED
@@
A prominent yet heuristic circumvention of this degeneracy is the Positive Degree-Day (PDD) scaling [5], which remains a standard parameterization in Earth System Models for projecting cryospheric responses across diverse timescales [8, 9, 7]. To estimate melt rates, this empirical approach assumes that annual cumulative mass loss is proportional to the time-integrated daily air temperature above the freezing point [1], effectively treating air temperature as a proxy for the energy input histories. While the PDD method exhibits remarkable predictive skill under relatively stable climatic regimes [14], its lack of a first-principles derivation from the surface energy balance introduces structural uncertainties. This limitation becomes particularly acute under non-stationary climate conditions, where the historical coupling between air temperature and the surface energy budget may diverge, potentially rendering existing parameterizations unreliable for future projections [10]. A fundamental gap therefore persists between the rigorous physics of energy-balance formulations and the practical efficacy of temperature-index schemes.
@@
Geometrically, since $\theta_{\mathrm{obs}}(t)=\theta_{f}$ throughout the melt season, the left-hand side of Eq. (6) corresponds to the $L^{1}$ distance between the latent and observed temperature trajectories. In one dimension, this distance is equivalent to the Wasserstein-1 metric $W_{1}(\theta,\theta_{obs})$ (see Appendix Section A). So
$W_{1}(\theta,\theta_{obs})=\tau M.$ (7)
Phase change may therefore be interpreted as an optimal transport process that compresses latent thermal fluctuations onto the melting boundary, with the melt energy determining the associated transport cost.
@@
Using representative values for ice-sheet surfaces ($C_{\rm sen}\approx 22.5\,\mathrm{W\,m^{-2}\,K^{-1}}$, $\theta_{f}=273.15\,\mathrm{K}$, $\sigma=5.67\times 10^{-8}\,\mathrm{W\,m^{-2}\,K^{-4}}$, $\rho_{i}=917\,\mathrm{kg\,m^{-3}}$, and $L_{f}=3.34\times 10^{5}\,\mathrm{J\,kg^{-1}}$) [3, 13], we obtain $C_{\rm PDD}\approx 7.6\,\mathrm{mm\,d^{-1}\,{}^{\circ}C^{-1}}$, well within the empirically inferred range ($1.5-8.5\mathrm{mm\,d^{-1}\,{}^{\circ}C^{-1}}$ for Greenland)[12]. This agreement indicates that the relaxation timescale $\tau$ emerging from the latent framework is quantitatively consistent with known radiative and turbulent dissipation processes.
@@
with amplitude $A=10$. The function is chosen such that the positive forcing phase spans approximately two-thirds of the cycle to better display. For clarity, we adopt a normalized thermal inertia $\rho C_{s}=1$ and reconstruct the latent temperature $\theta(t)$ subject to the energetic constraints described above and minimum $\mathcal{R}$ variation (see Appendix Section B).
@@ Acknowledgements.
Appendix A Latent temperature and the 1-Wasserstein distance
@@
$\int_{0}^{1}\bigl[\theta(t)-\theta_{\mathrm{obs}}(t)\bigr]\,dt=\int_{0}^{1}\left|F_{\theta}^{-1}(p)-F_{\theta_{\mathrm{obs}}}^{-1}(p)\right|\,dp.$ (14)
@@
Appendix B Thermodynamic Origin of Degree-Day Scaling in Phase-Change Systems
B.1 Governing Equation and General Solution
@@
$\rho C_{s}\frac{d\theta}{dt}+\frac{1}{\tau}(\theta-\theta_{f})=Q(t)-\mathcal{R}(t),$ (15)
@@
Equation (15) admits the formal solution
$\theta(t)=\theta_{f}+e^{-\lambda(t-t_{s})}\int_{t_{s}}^{t}\frac{Q(s)-\mathcal{R}(s)}{\rho C_{s}}e^{\lambda(s-t_{s})}\,ds,$ (16)
@@ B.2 Admissibility Constraints
$\int_{t_{s}}^{t_{e}}\mathcal{R}(t)\,dt=0.$ (17)
Second, closure of the latent trajectory at the end of the melt season, $\theta(t_{e})=\theta_{f}$, applied to Eq. (16), yields the weighted constraint
$\int_{t_{s}}^{t_{e}}\mathcal{R}(t)e^{\lambda t}\,dt=\int_{t_{s}}^{t_{e}}Q(t)e^{\lambda t}\,dt\equiv K.$ (18)
@@
B.3 Minimum Structural Perturbation Principle
@@
$J[\mathcal{R}]=\int_{t_{s}}^{t_{e}}\mathcal{R}(t)^{2}\,dt,$ (19)
subject to the constraints (17) and (18).
@@
$\mathcal{L}=\int_{t_{s}}^{t_{e}}\mathcal{R}^{2}\,dt+\mu_{1}\int_{t_{s}}^{t_{e}}\mathcal{R}\,dt+\mu_{2}\left(\int_{t_{s}}^{t_{e}}\mathcal{R}e^{\lambda t}\,dt-K\right).$ (20)
@@
$\frac{\delta\mathcal{L}}{\delta\mathcal{R}}=2\mathcal{R}(t)+\mu_{1}+\mu_{2}e^{\lambda t}=0,$ (21)
@@
$\mathcal{R}^{*}(t)=\alpha+\beta e^{\lambda t},$ (22)
@@
B.4 Determination of Coefficients
Substituting Eq. (22) into the two admissibility constraints yields the linear system
[…]
where $K$ is defined by Eq. (18). All integrals are analytic, ensuring a smooth and numerically stable solution for $\mathcal{R}^{*}(t)$ and $\theta(t)$ for any prescribed $\tau$.
B.5 Selection of the Dissipation Timescale
@@ B.6 Implementation
    Solve Eq. (23) for $\alpha$ and $\beta$.
@@
    Reconstruct $\theta(t)$ from Eq. (16) using $\mathcal{R}^{*}(t)$.